the difference between two numbers is 16. three times the larger number is seven times the smaller
Answer by checkley77(12844)

X-Y=16
X=16+Y
3X=7Y REPLACE X BY (16+Y)
3(16+Y)=7Y
48+3Y=7Y
3Y-7Y=-48
-4Y=-48
Y=-48/-4
Y=12 ANS. FOR THE SMALLER NUMBER.
X-12=16
X=16+12
X=28 ANS. FOR THE LARGER NUMBER.
PROOF:
3*28=7*12
84=84